The average monthly income of a person for first eight months of a year exceeds that for first seven months of the year by 25. If he earna 1275 in eight month of the year find his average monthly income for the first eight months of the year
Answer by Boreal(15235)

average = total/n
let average for 7 months =x
then x=total/7
when there are 8 months, the average is increased by 25 and the total increased by 1275, because that is what he earned.
so (x+25)=(total+1275)/8
7x=total; 8x=(8/7) total
8x+200=total +1275
8x=total+1075=(8/7) total
1075=(1/7) total, because (8/7) total-1 total =(1/7) total
total=$7525, and that is for 7 months
average for 7 months is $1075/month
total for 8 months is $8800, adding $1275.
average for 8 months is $8800/8=$1100, and that is $25 more than the 7 month average.
The answer is $1100 average monthly income for the first 8 months of the year